OpenStudy (anonymous):

Simplify by factoring. Assume that all expressions under radicals represent non-negative numbers. ^3√56*t^8

OpenStudy (saifoo.khan):

please write that clearly.

OpenStudy (anonymous):

\[3\sqrt{56\times t ^{8}?}\]

OpenStudy (anonymous):

\[\LaTeX\] \[6t^4\sqrt{14}\]